OpenStudy (anonymous):

Find all values of x on the interval 0<=x<=2Pi for which the graph of y=cos(x^2-1) has a horizontal tangent. Find all values of x on the interval 0<=x<=2Pi for which the graph of y=cos(x^2-1) has a horizontal tangent. @Mathematics

OpenStudy (anonymous):

i got this far y'=-2Xsin(x^2-1)

OpenStudy (amistre64):

good, and when does that equal zero? which is the slope of a horizontal line

OpenStudy (amistre64):

-2x = 0 sin(x^2-1) = 0 both are valid for solving

OpenStudy (anonymous):

i didnt know i can separate -2x from sin =) so is the answer just x=0,x=1,x=-1?

OpenStudy (amistre64):

:) its not so much as separating as seeing that if either of those goes to zero, the value of the whole goes to zero. -2x(0) = 0 0 sin(x^2-1) = 0 but yeah, 0 and -1 and 1 all look good

OpenStudy (amistre64):

we got a zero for sin(pi) as well

OpenStudy (amistre64):

-1 doesnt count since its not between 0 and 2pi

OpenStudy (anonymous):

i can really tell interval 0<=x<=2Pi

OpenStudy (anonymous):

can't

OpenStudy (amistre64):



OpenStudy (amistre64):

since x=-1 is not in that interval, we scrap it

OpenStudy (amistre64):

x^2 - 1 = pi is an option

OpenStudy (anonymous):

how do you know thats an option?

OpenStudy (amistre64):

becasue sin(pi) = 0

OpenStudy (anonymous):

thanks a lot